Solve using the Quadratic Formumla:
2y^2-(y+2)(y-3)=12
Please and Thanks!

Answer by TimothyLamb(4379)

2y^2 - (y + 2)(y - 3) = 12
2y^2 - (y^2 - y - 6) = 12
2y^2 - y^2 + y + 6 = 12
y^2 + y - 6 = 0
---
the above quadratic equation is in standard form, with a=1, b=1, and c=-6
---
to solve the quadratic equation, plug this:
1 1 -6
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
Answer 1:
the vertex of the quadratic equation is a minimum point at: ( -0.5, -6.25 )
---
Answer 2:
the quadratic equation has two real roots:
root1 = 2
root2 = -3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php